# Palpation Alters Auditory Pain Expressions with Gender-Specific Variations in Robopatients

### v1

@@ Abstract
Diagnostic errors remain a major cause of preventable deaths, particularly in resource-limited regions. Medical training simulators, including robopatients, play a vital role in reducing these errors by mimicking real patients for procedural training such as palpation. However, generating multimodal feedback, especially auditory pain expressions, remains challenging due to the complex relationship between palpation behavior and sound. The high-dimensional nature of pain sounds makes exploration challenging with conventional methods. This study introduces a novel experimental paradigm for pain expressivity in robopatients where they dynamically generate auditory pain expressions in response to palpation force, by co-optimizing human feedback using machine learning. Using Proximal Policy Optimization (PPO), a reinforcement learning (RL) technique optimized for continuous adaptation, our robot iteratively refines pain sounds based on real-time human feedback. This robot initializes randomized pain responses to palpation forces, and the RL agent learns to adjust these sounds to align with human preferences. The results demonstrated that the system adapts to an individual’s palpation forces and sound preferences and captures a broad spectrum of pain intensity, from mild discomfort to acute distress, through RL-guided exploration of the auditory pain space. The study further showed that pain sound perception exhibits saturation at lower forces with gender-specific thresholds. These findings highlight the system’s potential to enhance abdominal palpation training by offering a controllable and immersive simulation platform.
@@ I Introduction
Studying the multidimensional space of human pain requires extensive human studies and the analysis of large datasets. Conducting these studies under diverse conditions and collecting varied data is a time-intensive process, posing significant challenges in pain research. Machine learning algorithms can significantly enhance the design of human studies by adapting in real-time to participant behaviour [13], which can reduce the time taken to conduct these studies. This ensures that relevant data are collected efficiently, facilitates more accurate sampling, and reduces trial times by continuously refining protocols based on intermediate results.
This study introduces a novel framework for human-in-the-loop learning by incorporating Proximal Policy Optimisation to dynamically generate adaptive pain expressions in a robopatient during abdominal palpation. We used robopatient [7], a robotic platform that dynamically generates facial expressions and sounds in response to palpation force input from human participants during abdominal palpation. In contrast to traditional studies that collect data separately for post-analysis, this experiment follows a closed-loop approach, integrating real-time human feedback into the learning process, allowing the robot to dynamically adjust its auditory pain expressions, creating a more adaptive and personalised experience based on individual palpation. The concept of robopatient in this context is illustrated in Fig. 1. The results of the experiment demonstrated the variability of the relationship between palpation and auditory pain expression. Hence suggesting the complexity and multidimensional nature of pain expressions, as well as the possibility that this relationship could be personalised.
@@ II Related Work
The integration of human and ML approaches to generate adaptive auditory pain expressions in human physical twins intersects several research areas including, Artificial Intelligence (AI)-Driven robopatients, Multimodal Human-Robot Interaction (HRI) and objective pain assessment using AI. Robopatients use AI to simulate, predict, and optimize real-world processes such as simulating patient-specific scenarios to optimise treatment plans [14]. Advancements in multimodal HRI enable communication between humans and robots using various modalities, including visual, auditory and haptic information which facilitates more natural and intuitive interactions [15], which are crucial for applications involving emotional and pain recognition. Recent studies on objective pain assessment using AI have explored the use of machine learning models for pain recognition by analysing physiological signals. For instance, research utilising electrodermal activity (EDA) signals has applied mixup data augmentation and support vector machines to distinguish pain levels, achieving notable accuracy improvements [16]. These areas collectively contribute to the generating adaptive auditory pain expressions, enhancing the realism of robopatient.
@@ II-A Pain
Pain perception and expressions vary depending on various backgrounds such as gender [17] and ethnicity [18]. Infants experiencing a higher intensity of pain produce more irregular cries [19] with higher amplitude [20, 21], larger fundamental frequency variation [22, 23], smaller amplitude variation, and longer duration [22, 24]. Hence, the time scales of pain sound is also important to consider, for instance higher pitch and louder pain cries are associated with a sense of urgency [22, 25].
Individual experience of pain can significantly differ even for the same sensory input and development of combined psychophysical or psychological models that optimise pain treatment [26]. Hence it makes modelling the pain sounds or vocal pain expressions challenging [27, 28]. This background shows that generating an accurate model to represent subjective pain in someone is an extremely challenging task, which requires extensive study of the human perception regarding pain expressions.
@@
Figure 2: Physical robot and the experiment procedure A) Robopatient setup: A scenario where the participant is palpating on robopatient is shown. B) Data flow within the robopatient setup during the experiment. The interaction between the robot and the user starts as the user palpates on the abdominal phantom (Steps 1,2). The force applied by the user has to reach a certain level indicated by a progress bar next to the robot’s face (3). The progress bar is initially green and turns red when the required minimum amount of force is applied. Based on the palpation force, the robot generates random facial expressions and sounds independent to each other (4). Then the user listens to the pain sounds generated by the robot for the palpation force (5) and provides feedback on whether they agree with the robot’s mapping of pain sounds to palpation (6). This feedback updates the policy used by the PPO to generate future responses (7). This completes the one interaction cycle between the user and robot. Then robopatient learns the appropriate combinations of pitch and amplitude for pain sounds and optimises its response to maximise reward.
II-B Human-ML co-optimisation of emotion generation
@@
Learning algorithms can accommodate the multidimensionality of emotions and reduce bias, especially in the pain domain which is not well established yet in psychophysics. ML has proven effective skills in outperforming humans in analytical tasks such as bibliometry [31], emotion recognition and human studies [13]. By employing genetic algorithms, the study in [13] efficiently explored the high-dimensional space of facial expressions, demonstrating the importance of personalised approaches in understanding human emotions and improving emotion recognition systems. This study encourages moving away from stereotypical representations of emotions in research and towards approaches that capture the diversity and richness of emotional expression. In addition, prior knowledge improves our understanding of a situation, and prior experience accelerates our ability to adjust to new situations [32]. Hence AI for adaptive experiment design has been widely discussed in design optimisation [33], managing data and knowledge [34] in scientific discovery, identifying informative training samples [35], benchmarking datasets and democratising science.
@@
Human-ML co-optimisation involves combining human intuition with algorithmic optimization to enhance a robot’s performance. Previous studies have used Batch Bayesian Optimization to refine robot actions through human guidance, improving task efficiency [36] and compared human designers with model-free optimisation in a hopping robot [37], highlighting how co-optimisation can improve human creativity alongside ML.
@@
Current RL algorithms allow using small, high-quality datasets instead of large, noisy data. PPO is particularly well-suited for studying emotions and human behaviour due to its efficiency in handling smaller, high-quality datasets, making it ideal for real-time, interactive environments. Unlike other RL methods that may require vast amounts of noisy data, PPO proved to ensure stable policy and gradient updates [38]. Its scalability in high-dimensional spaces, including complex domains such as emotions and human behaviour, further underscores its applicability to our study, where the system must adapt to dynamic and nuanced interactions between participants and the robopatient [39]. Hence, we suggest a novel, hybrid approach for planning adaptive human experiments which typically are time-intensive and require large populations to derive useful conclusions. Pattern recognition and analytical skills that learning algorithms possess could make a paradigm shift in how we conduct traditional human studies and how collaboration between humans and robots can accelerate learning.
@@ III METHODS
The experimental setup, shown in Fig. 2A, involves users palpating a soft phantom laid on a rigid platform with four force sensors, which translates the palpation force into corresponding pain expressions produced by the system (robopatient). Robopatient is equipped with this force sensor platform on which a soft phantom is placed. This soft phantom acts as the abdomen. When users palpate on this soft phantom, the force sensor platform measures the palpation force as the input and generates appropriate auditory pain expressions as outputs in real-time. We combined this output with an appropriate facial expression, which is then projected onto the morphface: robopatient’s face, to make the scenario more interactive and realistic [40]. The user listens to the pain sound generated by the robot with the headset. As the next step, the user compares his/her palpation action and the robot’s pain expressions. The user then rates the congruence of these responses on a Likert scale with binary choices: Agree and Disagree. These inputs and feedback are fed to the PPO at each attempt to learn the preferences of the user. Over time, the PPO optimises pain sounds to maximise participant feedback scores, thereby improving the alignment between input force and the robot’s responses to pain.
@@ III-B Robopatient
Robopatient consists of the following hardware discussed in detail. The force sensor platform was made of four single-axis 20kg load cells (Haljia Inc) mounted on the four corners of a generic square shaped (250$\times$250 mm) acrylic platform, topped with a Silicone phantom (200$\times$200$\times$30 mm, Ecoflex 00-30). The distance between two adjacent load cells is 246 mm. A National Instruments DAQ (USB-6212) with MATLAB 2024a is used to acquire the signals from the load cells at a sampling rate of 1000 Hz. The sum of the force values from each load cell is used to calculate the total palpation force $F_{tp}$. To minimise the noise in these calculations, any load cell value where the total palpation force is less than 0.5N has been neglected based on the observations from pilot experiments. $F_{tp}$ can be calculated as follows.
@@ III-C Palpation to Pain Intensity Mapping
$Amplitude\ Level=[1,0.3,0.1,0.037]^{T}\vspace{-1.5em}$ (2)
$Pitch\ Level=[0.7,0.9,1.1,1.3]^{T}\vspace{-1.5em}$ (3)
$Force=[5,10,15,20]^{T}\vspace{-1.5em}$ (4)
@@
III-D Dynamic Interaction Optimisation between the User and Robopatient
We utilise PPO algorithm, an RL method known for its stability and efficiency in real-time environments [38, 41]. PPO is particularly well-suited for our application as it can handle continuous learning while maintaining the stability of policy updates through a clipped surrogate objective function. This mechanism prevents large deviations in policy updates, reducing the risk of destabilizing learning and ensuring consistent adaptation over time.
While PPO is a widely used RL method, it is important to consider alternative RL approaches that could potentially be applied to our problem. Several RL methods can be use for continuous action spaces, each with its own advantages and limitations. Soft Actor Critic(SAC) [42] is known for sample efficiency and robust performance in high-dimensional tasks but requires significant computational resources, leading to less suitable for real-world applications. Deep Deterministic Policy Gradient (DDPG) [43] and its advanced version Twin Delayed DDPG (TD3) [44] are widely used in continuous control problems but are highly sensitive to hyperparameter tuning and often suffer from training instability. Given that our application involves continuous user interaction and real-time adaptation, PPO offers the best trade-off between training efficiency and policy stability, making it the ideal choice for this study.
@@ III-D Dynamic Interaction Optimisation between the User and Robopatient
The PPO algorithm maintains two neural networks: a policy network, which selects actions (displaying appropriate pitch and amplitude) based on palpation force, and a value network, which estimates future rewards. The policy is iteratively updated through mini-batch gradient descent, optimising the clipped objective to balance learning stability and performance.
@@
The workflow of PPO used in this study is illustrated in Fig.4. The algorithm starts by initialising its policy and value network as it sets up the interaction environment involving the robot and the user. For each interaction, the algorithm observes the current palpation data and selects a sound combination (pitch and amplitude) using the policy network. The selected action is then presented to the user, who provides binary feedback (Agree or Disagree), which is taken as a reward. The system updates its state based on the feedback and stores transitions (state, action, reward, next state) for policy and value updates. PPO optimises the policy network, while the value network is updated to improve future rewards. Then it checks for a termination condition after each episode, and if met, the episode ends, displaying the most successful action-input mapping. If the condition is not met, the system continues learning; otherwise, the environment resets and starts a new episode.
Figure 4: Process flow diagram of the PPO algorithm

### v2

@@ Abstract
Diagnostic errors remain a major cause of preventable deaths, particularly in resource-limited regions. Medical training simulators, including robopatients, play a vital role in reducing these errors by mimicking real patients for procedural training such as palpation. However, generating multimodal feedback, especially auditory pain expressions, remains challenging due to the complex relationship between palpation behavior and pain sound. The high-dimensional nature of pain sounds makes exploration challenging with conventional methods. This study introduces a novel experimental paradigm for pain expressivity in robopatients where they dynamically generate auditory pain expressions in response to palpation force, by co-optimizing realtime human feedback using machine learning. Using Proximal Policy Optimization (PPO), a reinforcement learning (RL) technique optimized for continuous adaptation, our robot iteratively refines pain sounds based on real-time human feedback. This robot initializes randomized pain responses to palpation forces, and the RL agent learns to adjust these sounds to align with human preferences. The results demonstrated that the system adapts to an individual’s palpation forces and sound preferences and captures a broad spectrum of pain intensities, from mild discomfort to acute distress. The study further showed that pain sound perception exhibits saturation at lower forces with gender-specific thresholds. Thus, this study presents an early attempt to use human-in-the-loop reinforcement learning to co-optimize haptic input and auditory pain expression. These findings highlight the system’s potential to enhance abdominal palpation training by offering a controllable and immersive simulation platforms.
@@ I Introduction
Studying the multidimensional space of human pain requires extensive human studies and the analysis of large datasets. Conducting these studies under diverse conditions and collecting varied data is a time-intensive process, posing significant challenges in pain research. Machine learning algorithms can significantly enhance the design of human studies by adapting in real-time to participant behavior [1], which can reduce the time taken to conduct these studies. This ensures that relevant data are collected efficiently, facilitates more accurate sampling, and reduces trial times by continuously refining protocols based on intermediate results.
This study introduces a novel framework for human-in-the-loop learning by incorporating Proximal Policy optimization to dynamically generate adaptive pain expressions in a robopatient during abdominal palpation. This work highlights early trends in how auditory expressions of pain can be shaped through interactive learning with human feedback. We used robopatient [33], a robotic platform that dynamically generates facial expressions and sounds in response to palpation force input from human participants during abdominal palpation. In contrast to traditional studies that collect data separately for post-analysis, this experiment follows a closed-loop approach, integrating real-time human feedback into the learning process, allowing the robot to dynamically adjust its auditory pain expressions, creating a more adaptive and personalised experience based on individual palpation. Hence, this contribution can be considered an early step toward making the planning and execution of human studies on high-dimensional phenomena such as pain less time-consuming and less exhausting, by incorporating co-optimization and real-time human feedback. The concept of robopatient in this context is illustrated in Fig. 1. The results of the experiment demonstrated the variability of the relationship between palpation and auditory pain expression. Hence suggesting the complexity and multidimensional nature of pain expressions, as well as the possibility that this relationship could be personalised.
@@ II Related Work
The integration of human and ML approaches to generate adaptive auditory pain expressions in human physical twins intersects several research areas including, Artificial Intelligence (AI)-Driven robopatients, Multimodal Human-Robot Interaction (HRI) and objective pain assessment using AI.
Robopatients use AI to simulate, predict, and optimize real-world processes such as simulating patient-specific scenarios to optimize treatment plans [31]. Advancements in multimodal HRI enable communication between humans and robots using various modalities, including visual, auditory and haptic information which facilitates more natural and intuitive interactions [43], which are crucial for applications involving emotional and pain recognition. Recent studies on objective pain assessment using AI have explored the use of machine learning models for pain recognition by analysing physiological signals. For instance, research utilising electrodermal activity (EDA) signals has applied mixup data augmentation and support vector machines to distinguish pain levels, achieving notable accuracy improvements [40]. These areas collectively contribute to the generating adaptive auditory pain expressions, enhancing the realism of robopatient.
@@ II-A Pain
Pain perception and expressions vary depending on various backgrounds such as gender [35] and ethnicity [34]. Infants experiencing a higher intensity of pain produce more irregular cries [45] with higher amplitude [28, 12], larger fundamental frequency variation [32, 23], smaller amplitude variation, and longer duration [32, 21]. Hence, the time scales of pain sound are also important to consider, for instance higher pitch and louder pain cries are associated with a sense of urgency [32, 7]. However, when synthesizing pain artificially, design choices must ensure that the output remains within the human capacity to reliably interpret and respond—typically limited to no more than four simultaneous variables in affective or perceptual tasks [2, 16].
Individual experience of pain can significantly differ even for the same sensory input and development of combined psychophysical or psychological models that optimize pain treatment [5]. Hence it makes modelling the pain sounds or vocal pain expressions challenging [27, 18]. This background shows that generating an accurate model to represent subjective pain in someone is an extremely challenging task, which requires extensive study of the human perception regarding pain expressions.
@@
Figure 2: Physical robot and the experiment procedure A) Robopatient setup: A scenario where the participant is palpating on robopatient is shown. B) Data flow within the robopatient setup during the experiment. The interaction between the robot and the user starts as the user palpates on the abdominal phantom (Steps 1, 2). The force applied by the user has to reach a certain level indicated by a progress bar next to the robot’s face (step 3). The progress bar is initially green and turns red when the required minimum amount of force is applied. Based on the palpation force, the robot generates random facial expressions and sounds independent to each other (step 4). Then the user listens to the pain sounds generated by the robot for the palpation force (step 5) and provides feedback on whether they agree with the robot’s mapping of pain sounds to palpation (step 6). This feedback updates the policy used by the PPO to generate future responses (step 7). This completes the one interaction cycle between the user and robot. Then robopatient learns the appropriate combinations of pitch and amplitude for pain sounds and optimizes its response to maximise reward.
II-B Human-ML co-optimization for emotion generation
@@
Learning algorithms can accommodate the multidimensionality of emotions and reduce bias, especially in the pain domain which is not well established yet in psychophysics. ML has proven effective skills in outperforming humans in analytical tasks such as bibliometry [14], emotion recognition and human studies [1]. By employing genetic algorithms, the study in [1] efficiently explored the high-dimensional space of facial expressions, demonstrating the importance of personalised approaches in understanding human emotions and improving emotion recognition systems. This study encourages moving away from stereotypical representations of emotions in research and towards approaches that capture the diversity and richness of emotional expression. In addition, prior knowledge improves our understanding of a situation, and prior experience accelerates our ability to adjust to new situations [48]. Hence AI for adaptive experiment design has been widely discussed in design optimization [4], managing data and knowledge [50] in scientific discovery, identifying informative training samples [46], benchmarking datasets and democratizing science.
@@
Human-ML co-optimization involves combining human intuition with algorithmic optimization to enhance a robot’s performance. Previous studies have used Batch Bayesian Optimization to refine robot actions through human guidance, improving task efficiency [22] and compared human designers with model-free optimization in a hopping robot [36], highlighting how co-optimization can improve human creativity alongside ML.
@@
Current RL algorithms allow using small, high-quality datasets instead of large, noisy data. PPO is particularly well-suited for studying emotions and human behavior due to its efficiency in handling smaller, high-quality datasets, making it ideal for real-time, interactive environments. Unlike other RL methods that may require vast amounts of noisy data, PPO proved to ensure stable policy and gradient updates [38]. Its scalability in high-dimensional spaces, including complex domains such as emotions and human behavior, further underscores its applicability to our study, where the system must adapt to dynamic and nuanced interactions between participants and the robopatient [9]. Hence, we suggest a novel, hybrid approach for planning adaptive human experiments which typically are time-intensive and require large populations to derive useful conclusions. Pattern recognition and analytical skills that learning algorithms possess could make a paradigm shift in how we conduct traditional human studies and how collaboration between humans and robots can accelerate learning.
@@ III METHODS
The experimental setup, shown in Fig. 2A, involves users palpating a soft phantom laid on a rigid platform with four force sensors, which translates the palpation force into corresponding pain expressions produced by the system (robopatient). Robopatient is equipped with this force sensor platform on which a soft phantom is placed. This soft phantom acts as the abdomen. When users palpate on this soft phantom, the force sensor platform measures the palpation force as the input and generates appropriate auditory pain expressions as outputs in real-time. We combined this output with an appropriate facial expression, which is then projected onto the morphface: robopatient’s face, to make the scenario more interactive and realistic [6]. The user listens to the pain sound generated by the robot with the headset. As the next step, the user compares his/her palpation action and the robot’s pain expressions. The user then rates the congruence of these responses on a Likert scale with binary choices: Agree and Disagree. These inputs and feedback are fed to the PPO at each attempt to learn the preferences of the user. Over time, the PPO optimizes pain sounds to maximise participant feedback scores, thereby improving the alignment between input force and the robot’s responses to pain.
@@ III-B Robopatient
Robopatient consists of the following hardware discussed in detail. The force sensor platform was made of four single-axis 20kg load cells (Haljia Inc) mounted on the four corners of a generic square shaped ($\times250$ mm) acrylic platform, topped with a Silicone phantom ($\times200$$\times$30 mm, Ecoflex 00-30). The distance between two adjacent load cells is 246 mm. A National Instruments DAQ (USB-6212) with MATLAB 2024a is used to acquire the signals from the load cells at a sampling rate of 1000 Hz. The sum of the force values from each load cell is used to calculate the total palpation force $F_{tp}$. To minimise the noise in these calculations, any load cell value where the total palpation force is less than 0.5N has been neglected based on the observations from pilot experiments. $F_{tp}$ can be calculated as follows.
@@ III-C Palpation to Pain Intensity Mapping
$Amplitude\ Level=[1,0.3,0.1,0.037]^{T}\vskip-15.00002pt$ (2)
$Pitch\ Level=[0.7,0.9,1.1,1.3]^{T}\vskip-15.00002pt$ (3)
$Force=[5,10,15,20]^{T}\vskip-15.00002pt$ (4)
@@
III-D Dynamic Interaction optimization between the User and Robopatient
We utilise PPO algorithm, an RL method known for its stability and efficiency in real-time environments [38, 19]. PPO is particularly well-suited for our application as it can handle continuous learning while maintaining the stability of policy updates through a clipped surrogate objective function. This mechanism prevents large deviations in policy updates, reducing the risk of destabilizing learning and ensuring consistent adaptation over time. By embedding human evaluators in the learning loop, the robot’s sound responses adapt not only to the applied force but also to the subjective preferences and interpretations of the user. Our objective in this study is to explore the feasibility of using this human–machine learning co-optimization approach to generate adaptive, expressive auditory pain responses during palpation rather than aiming to identify the best-performing algorithm.
@@ III-D Dynamic Interaction optimization between the User and Robopatient
III-E PPO Algorithm
Compared to alternative RL approaches, PPO demonstrates critical advantages for our human-in-the-loop paradigms: Its sample efficiency minimizes trial burden on participants given sparse binary feedback; stable convergence via clipped objectives avoids policy updates during continuous interactions.
Soft Actor Critic(SAC) [15] offers sample efficiency but requires significant computational resources, making it impractical for real-time deployment. Deep Deterministic Policy Gradient (DDPG) [29] and its advanced version Twin Delayed DDPG (TD3) [11] suffer from hyperparameter sensitivity and training instability, which could disrupt human-robot interaction continuity.
Given our application involves continuous user interaction and real-time adaptation under constrained feedback, PPO provides the optimal balance of training efficiency, policy stability, and practical deployability.
The problem was formulated as an episodic Markov Decision Process (MDP), where the state at each time step is defined by the palpation force input, and the reward is based solely on the resulting response on the pain expression. Although human behavior introduces inherent stochasticity, each palpation episode is reset independently, and the user’s next input is not influenced by the robot’s action, thus satisfying the assumptions of an MDP.
As identified earlier, using a clipped objective function in the PPO prevents large updates that could destabilize the training process, ensuring that policy updates remain within a reasonable range and promoting stable learning. Additionally, PPO maintains and updates both the policy network and the value function network. The policy network determines the actions to take based on the current state, while the value function network estimates the expected return, aiding in the improvement of the policy’s decision-making process.
@@ III-E PPO Algorithm
The PPO algorithm maintains two neural networks: a policy network, which selects actions (selecting appropriate pitch and amplitude) based on the palpation force, and a value network, which estimates future rewards. The policy is iteratively updated through mini-batch gradient descent, optimizing the clipped objective to balance learning stability and performance. Concretely, we use separate actor/critic MLPs with two hidden layers of $64$ units and tanh activations, optimized with Adam (learning rate $3\times 110^{-4}$), minibatch size $64$, and $10$ epochs per update; these choices follow the common PPO practice [38, 19]. The pilot runs showed the fastest stable convergence, whereas training was slower at $1{\times}10^{-4}$ and oscillatory at $1{\times}10^{-3}$.
@@
The workflow of PPO used in this study is illustrated in Fig.4. The algorithm starts by initialising its policy and value network as it sets up the interaction environment involving the robot and the user. For each interaction, the algorithm observes the current palpation data and selects a sound combination (pitch and amplitude) using the policy network. The selected action is then presented to the user, who provides binary feedback (Agree or Disagree), which is taken as a reward. The system updates its state based on the feedback and stores transitions (state, action, reward, next state) for policy and value updates. PPO optimizes the policy network, while the value network is updated to improve future rewards. Then it checks for a termination condition after each episode, and if met, the episode ends, displaying the most successful action-input mapping. If the condition is not met, the system continues learning; otherwise, the environment resets and starts a new episode. We terminate when the last 10 iterations contain $\geq 8/10$ Agree responses. This $80\%$ threshold was selected in pilot sessions as the most stable stopping rule that avoids premature convergence while limiting participant fatigue.
A detailed workflow of applying PPO is summarized in the following steps:
  1. 1.
    Initialization: The system initializes its policy and value networks, which are neural networks responsible for deciding the appropriate facial expression and combinatorial sound parameters to respond with, based on the palpation force (input) from the participant. The environment, which consists of the robopatient and the participant, is also initialized.
  2. 2.
    Action Selection: For each interaction, the system observes the current input and uses the policy network to select an action, which is a combination of a facial expression and sounds parameters.
  3. 3.
    Execution and Feedback: The selected action is presented to the participant, who provides feedback on how appropriate they found the response. This feedback is a score from 1 to 4.
  4. 4.
    Reward and Observation Update: The system receives the feedback score as a reward and updates the current state based on the new force input. The transition (current state, action, reward, next state) is stored for policy and value updates.
  5. 5.
    Policy and Value Network Update: Using the collected transitions, the PPO algorithm computes the advantage estimates and updates the policy network by optimizing the clipped surrogate objective function. Simultaneously, the value network is updated to better predict the expected future rewards.
  6. 6.
    Termination Check: The system checks if the termination condition is met (currently set to the probability of achieving a feedback score of ‘strongly agree’ for ten consecutive interactions to $80\%$). If the condition is met, the episode ends, and the system displays the mapping between actions and force inputs that led to the highest participant satisfaction.
  7. 7.
    Reset and Continue: If the termination condition is not met, the system continues to the next iteration. If it is met, the environment resets, and a new episode begins.
Figure 4: Process flow diagram of the PPO algorithm.